# MindEye2: Shared-Subject Models Enable fMRI-To-Image With 1 Hour of Data

### v1

@@ 2.1 Shared-Subject Functional Alignment
Two strengths of this novel functional alignment procedure are in its simplicity and flexibility. Using a simple linear mapping for alignment can provide robust, generalizeable performance in low-sample, high-noise settings because simple mappings are less likely to overfit to noise. Also, unlike typical functional alignment approaches that require subjects to process a shared set of images (Haxby et al., 2011), our approach has the flexibility to work even when subjects are viewing entirely unique images in the training data. This is critical for the Natural Scenes Dataset, where 90% of the seen images are unique to the subject and the 10% that were seen across subjects are relegated to the test set. Further, this approach holds advantages for subsequent data collection of a new subject, where such data collection does not need to be restricted to showing a predefined set of images.
@@ 2.5 Model Inference
The final "refined" reconstructions come from combining the outputs from base SDXL with the pixel images output from the low-level submodule via simple weighted averaging (4:1 ratio). This weighted averaging step increases performance on low-level image metrics while minimally affecting reconstructions’ subjective appearance.
@@ 3 Results
We used the Natural Scenes Dataset (NSD) (Allen et al., 2022), a public fMRI dataset containing the brain responses of human participants viewing rich naturalistic stimuli from COCO (Lin et al., 2014). The dataset spans 8 subjects who were each scanned for 30-40 hours (30-40 separate scanning sessions), where each sesssion consisted of viewing 750 images for 3 seconds each. Images were seen 3 times each across the sessions and were unique to each subject, except for a select 1,000 images which were seen by all the subjects. We follow the standardized approach to train/test splits used by other NSD reconstruction papers (Takagi and Nishimoto, 2022; Ozcelik and VanRullen, 2023; Gu et al., 2023) which is to use the shared images seen by all the subjects as the test set. We follow the standard of evaluating model performance across low- and high-level image metrics averaged across the 4 subjects who completed all 40 scanning sessions. We averaged across same-image repetitions for the test set (1,000 test samples) but not the training set (30,000 training samples). For more information on NSD and data preprocessing see Appendix A.2.
Critically, models trained on a subset of data were selected in chronological order. That is, models trained from only $1$ hour’s worth of data come from using the subject’s first scanning session of 750 image presentations. This means our model must be able to generalize to test data collected from scanning sessions entirely held-out during training.
@@
First, we report performance of MindEye2 when training on the full NSD dataset. We quantitatively compare reconstructions across fMRI-to-image models in Table 1, demonstrating state-of-the-art MindEye2 performance across nearly all metrics. We compare to both the previous MindEye1 results as well as other fMRI-to-image approaches that were open-sourced such that we could replicate their pipelines using the recently updated NSD (which includes an additional 3 scanning sessions for every subject).
@@
We also report performance for MindEye2 fine-tuned with only 1 hour of data in the same Table 1. We qualitatively compare reconstructions side-by-side with models trained on only 1 hour’s worth of data in Figure 3, depicting improvements in reconstruction quality for MindEye2. We report more evaluations in the Appendix: see A.3 for MindEye2 results without pretraining, A.4 for evaluations with varying amounts of training data across all models, A.5 for single-subject evaluations, and A.10 for MindEye2 evaluations with varying selection of pretraining subjects. We also conducted a behavioral experiment with human raters which confirmed that humans subjectively prefer MindEye2 (1-hour) reconstructions to Brain Diffuser (1-hour) reconstructions (Appendix A.13).
@@ 4 Related Work
It is common for fMRI analyses to align subjects’ brains to a shared space for the purposes of increasing statistical power and/or assessing generality of scientific findings. Such alignment is difficult because structural and functional topography differs substantially across people (Talairach and Tournoux, 1990; Mazziotta et al., 2001). There are many approaches to functional alignment but typically they involve subjects experiencing shared stimuli and then using responses to these stimuli to learn an alignment mapping (Chen et al., 2015; Haxby et al., 2011; Huang et al., 2021; Nastase et al., 2019; Busch et al., 2021). While it is useful to conduct such experiments to identify sources of shared signal across subjects, it is also limiting in that new subjects would need to be scanned using the same experimental protocol. Other functional alignment approaches avoid such limitations by using self-supervised learning to identify an initial generalizable embedding space with outputs suitable for downstream tasks (Schneider et al., 2023; Chen et al., 2023a, b). Closest to our alignment approach are models that adopt both shared-subject and subject-specific mappings in their model architecture (Défossez et al., 2022; Benchetrit et al., 2023; Yang et al., 2023; Lane and Kiar, 2023).
@@ 5.1 Limitations
5.2 Broader Impacts
The present work demonstrates that it is now practical for patients to undergo a single MRI scanning session and produce enough data to perform high-quality reconstructions of their visual perception. Such image reconstructions from brain activity are expected to be systematically distorted due to factors including mental state, neurological conditions, etc. This could potentially enable novel clinical diagnosis and assessment approaches, including applications for improved locked-in (pseudocoma) patient communication (Monti et al., 2010) and brain-computer interfaces if adapted to real-time analysis (Wallace et al., 2022) or non-fMRI neuroimaging modalities. As technology continues to improve, we note it is important that brain data be carefully protected and companies collecting such data be transparent with their use.
@@ A.1 Author Contributions
PSS: project lead, drafted the initial manuscript and contributed to all parts of MindEye2 development. MT (core contributor): MindEye2 ablations, SDXL unCLIP vs. Versatile Diffusion comparisons, improved distributed training code, and experimented with approaches not used in the final model including training custom ControlNet and T2I adapters, using retrieval on COCO CLIP captions, and using diffusion priors to align fMRI to text embeddings. CKTV (core contributor): retrained and evaluated MindEye1 models, image captioning evaluations and writing, improved manuscript formatting, ROI-optimized stimuli experiments, and experimented with approaches not used in the final model including trying out different pretrained model embeddings, experimenting with T2I-Adapters and depth conditioning, experimenting with using past/future timepoints as additional conditioning, experimenting with blip2 for text prediction, and experimenting with behavioral embeddings. RK (core contributor): brain correlations, human preference experiments, recalculated metrics for 40-hour setting Ozcelik and VanRullen (2023) and Takagi and Nishimoto (2023) results, evaluations with varying amounts of training data across all models, assistance with data normalization, significant contributions to manuscript writing. TC: UMAP visualizations, improved the design for Figure 1, and experimented with approaches not used in the final model including using past/future timepoints as additional conditioning and using flattened voxels in MNI space instead of native space. AN: helped with ablations and experimented with replacing soft CLIP loss with soft SigLIP loss (Zhai et al., 2023) (not used in final model). CS: FAISS retrieval with MS-COCO (Appendix A.8) and experimented with approaches not used in the final model including experimenting with using past/future timepoints as additional conditioning, experimenting with blip2 for text prediction, and experimenting with behavioral embeddings. JX: helped with ablations, manuscript revisions and table formatting, experimented with approaches not used in the final model including experimenting with blip2 for text prediction, experimenting with behavioral embeddings, and improving model architecture. TN: assisted with human preference experiments. KN: oversaw the project, manuscript revisions and framing. TMA: oversaw the project, manuscript revisions and framing, helped keep project on-track thorugh MedARC and Stability AI communication.
@@ A.9 Reconstruction Evaluations: Additional Information
A.10 Pretraining with Less Subjects
@@
A.11 UMAP Dimensionality Reduction
A.12 ROI-Optimized Stimuli
Here we try to visualize the functional organization of the brain by feeding synthetic brain activity through pretrained MindEye2. Inspired by the ROI-optimal analyses of Ozcelik and VanRullen (2023), we utilized four ROIs derived from population receptive field (pRF) experiments and four ROIs derived from functional localization (fLoc) experiments. These pRF and fLoc experiments were provided by the NSD dataset. The ROIs are as follows (region names following the terminology adopted in Allen et al. (2021)): V1 is the concatenation of V1 ventral (V1v) and V1 dorsal (V1d), and similarly for V2 and V3; V4 is the human V4 (hV4); the Face-ROI consists of the union of OFA, FFA-1, FFA-2, mTL-faces, and aTL-faces; the Word-ROI consists of OWFA, VWFA-1, VWFA-2, mfs-words, and mTL-words; the Place-ROI consists of OPA, PPA, and RSC; and the Body-ROI consists of EBA, FBA-1, FBA-2, and mTL-bodies.
@@ A.12 ROI-Optimized Stimuli
Figure 10: Unrefined reconstructions and decoded captions from syntetic fMRI activity by simulating full and specific ROI activation.
@@
Such results demonstrate the potential to directly visualize preferential stimuli for any desired region of interest; further functional specialization exploration could be performed using more sophisticated methods (c.f., Sarch et al. (2023); Luo et al. (2023a, b)).
A.13 Human Preference Experiments
@@
A.13.1 MindEye2 vs. MindEye2 (random)
@@
A.13.2 MindEye2 (refined) vs. MindEye2 (unrefined)
@@
A.13.3 MindEye2 (1-hour) vs. Brain Diffuser (1-hour)

### v2

@@ 2.1 Shared-Subject Functional Alignment
Two strengths of this novel functional alignment procedure are in its simplicity and flexibility. Using a simple linear mapping for alignment can provide robust, generalizeable performance in low-sample, high-noise settings because simple mappings are less likely to overfit to noise. Also, unlike typical functional alignment approaches that require subjects to experience the same shared set of images (Haxby et al., 2011), our approach has the flexibility to work even when subjects are viewing entirely unique images in the training data. This is critical for the Natural Scenes Dataset, where 90% of the seen images are unique to the subject and the 10% that were seen across subjects are relegated to the test set. Further, this approach holds advantages for data collection of a new subject, where such data collection does not need to be restricted to showing a predefined set of images. This approach is also relevant for application on small target datasets where other pre-training data is available.
@@ 2.5 Model Inference
The final "refined" reconstructions come from combining the outputs from base SDXL with the pixel images output from the low-level submodule via simple weighted averaging (4:1 ratio). This weighted averaging step increases performance on low-level image metrics while minimally affecting reconstructions’ subjective appearance, but is overall not a critical component to MindEye2 and can be entirely discarded without a noticeable drop in reconstruction quality. Replacing this weighted averaging approach with conditioning the diffusion model using VAE embeddings as done in MindEye1 resulted in worsened performance, likely due to the OpenCLIP embeddings already doing a good job at retaining low-level image information.
@@ 3 Results
We used the Natural Scenes Dataset (NSD) (Allen et al., 2022), a public fMRI dataset containing the brain responses of human participants viewing rich naturalistic stimuli from COCO (Lin et al., 2014). The dataset spans 8 subjects who were each scanned for 30-40 hours (30-40 separate scanning sessions), where each sesssion consisted of viewing 750 images for 3 seconds each. Images were seen 3 times each across the sessions and were unique to each subject, except for a select 1,000 images which were seen by all the subjects. We follow the standardized approach to train/test splits used by other NSD reconstruction papers (Takagi and Nishimoto, 2022; Ozcelik and VanRullen, 2023; Gu et al., 2023) which is to use the shared images seen by all the subjects as the test set. We follow the standard of evaluating model performance across low- and high-level image metrics averaged across the 4 subjects who completed all 40 scanning sessions. We averaged across same-image repetitions for the test set (1,000 test samples) but not the training set (30,000 training samples). The inputs to the model during training and fine-tuning are always single-trial, individual (non-pooled) brain-image paired samples. For more information on NSD and data preprocessing see Appendix A.2.
Critically, models trained/fine-tuned on a subset of data were selected in chronological order. That is, models fine-tuned from only 1 hour’s worth of data come from using the subject’s first scanning session of 750 image presentations. These are individual samples not pooled across image repeats. This means our model must be able to generalize to test data collected from scanning sessions entirely held-out during training/fine-tuning and which involve reconstructing images never presented to the model during training or fine-tuning.
@@
First, we report performance of MindEye2 when training on the full NSD dataset. We quantitatively compare reconstructions across fMRI-to-image models in Table 1, demonstrating state-of-the-art MindEye2 performance across nearly all metrics. We compare to both the previous MindEye1 results as well as other fMRI-to-image approaches that were open-sourced such that we could replicate their pipelines using the recently updated NSD (which includes an additional 3 scanning sessions for every subject). For exhaustive figures depicting all MindEye2 reconstructions in the test set, see the figs folder in our GitHub repo.
@@
We also report performance for MindEye2 fine-tuned with only 1 hour of data in the same Table 1. We qualitatively compare reconstructions side-by-side with models trained on only 1 hour’s worth of data in Figure 3, depicting improvements in reconstruction quality for MindEye2. We report more evaluations in the Appendix: see A.3 for MindEye2 results without pretraining, A.4 for evaluations with varying amounts of training data across all models, A.5 for single-subject evaluations, A.11 for MindEye2 evaluations with varying selection of pretraining subjects, and A.14 for visualization of the functional preferences of various brain regions of interest along the visual hierarchy (Serre et al., 2005) for each subject. We also conducted a behavioral experiment with human raters which confirmed that humans subjectively prefer MindEye2 (1-hour) reconstructions to Brain Diffuser (1-hour) reconstructions (Appendix A.15).
@@ 4 Related Work
It is common for fMRI analyses to align subjects’ brains to a shared space for the purposes of increasing statistical power and/or assessing generality of scientific findings. Such alignment is difficult because structural and functional topography differs substantially across people (Talairach and Tournoux, 1990; Mazziotta et al., 2001). Anatomical alignment, where brains are physically warped into a predefined brain template, is imperfect and can potentially distort functional alignment (Fischl et al., 1999; Sabuncu et al., 2010; Conroy et al., 2009; Thual et al., 2022). Functional alignment can ignore anatomical structure and focus specifically on finding shared patterns of brain activity. There are many approaches to functional alignment but typically they involve subjects experiencing shared stimuli and then using responses to these stimuli to learn an alignment mapping (Chen et al., 2015; Haxby et al., 2011, 2020; Huang et al., 2021; Nastase et al., 2019; Busch et al., 2021). While it is useful to conduct such experiments to identify sources of shared signal across subjects, it is also limiting in that new subjects would need to be scanned seeing the same images. Other functional alignment approaches avoid such limitations by using self-supervised learning to identify an initial generalizable embedding space with outputs suitable for downstream tasks (Schneider et al., 2023; Chen et al., 2023a, b). Closest to our alignment approach are models that adopt both shared-subject and subject-specific mappings in their model architecture (Défossez et al., 2022; Benchetrit et al., 2023; Yang et al., 2023; Lane and Kiar, 2023).
@@ 5.1 Limitations
5.2 Impact Statement
The present work demonstrates that it is now practical for patients to undergo a single MRI scanning session and produce enough data to perform high-quality reconstructions of their visual perception. Such image reconstructions from brain activity are expected to be systematically distorted due to factors including mental state, neurological conditions, etc. This could potentially enable novel clinical diagnosis and assessment approaches, including applications for improved locked-in (pseudocoma) patient communication (Monti et al., 2010) and brain-computer interfaces if adapted to real-time analysis (Wallace et al., 2022) or non-fMRI neuroimaging modalities. Future work could potentially generalize reconstruction models from perception to mental imagery without training a new model (Stokes et al., 2009; Goebel et al., 2022; Naselaris et al., 2015; Reddy et al., 2010). As technology continues to improve, we note it is important that brain data be carefully protected and companies collecting such data be transparent with their use.
6 Author Contributions
For detailed author contributions see Appendix A.1.
@@ A.1 Author Contributions
PSS: project lead, drafted the initial manuscript and contributed to all parts of MindEye2 development. MT (core contributor): MindEye2 ablations, SDXL unCLIP vs. Versatile Diffusion comparisons, improved distributed training code, and experimented with approaches not used in the final model including training custom ControlNet and T2I adapters, using retrieval on COCO CLIP captions, and using diffusion priors to align fMRI to text embeddings. CKTV (core contributor): retrained and evaluated MindEye1 models, image captioning evaluations and writing, improved manuscript formatting, ROI-optimized stimuli experiments, and experimented with approaches not used in the final model including trying out different pretrained model embeddings, experimenting with T2I-Adapters and depth conditioning, experimenting with using past/future timepoints as additional conditioning, experimenting with blip2 (Li et al., 2023) for text prediction, and experimenting with behavioral embeddings. RK (core contributor): brain correlations, human preference experiments, recalculated metrics for 40-hour setting Ozcelik and VanRullen (2023) and Takagi and Nishimoto (2023) results, evaluations with varying amounts of training data across all models, assistance with data normalization, significant contributions to manuscript writing. TC: UMAP visualizations, improved the design for Figure 1, and experimented with approaches not used in the final model including using past/future timepoints as additional conditioning and using flattened voxels in MNI space instead of native space. AN: helped with ablations and experimented with replacing soft CLIP loss with soft SigLIP loss (Zhai et al., 2023) (not used in final model). CS: FAISS retrieval with MS-COCO (Appendix A.8) and experimented with approaches not used in the final model including experimenting with using past/future timepoints as additional conditioning, experimenting with blip2 for text prediction, and experimenting with behavioral embeddings. JX: helped with ablations, manuscript revisions and table formatting, experimented with approaches not used in the final model including experimenting with blip2 for text prediction, experimenting with behavioral embeddings, and improving model architecture. TN: assisted with human preference experiments. KN: oversaw the project, manuscript revisions and framing. TMA: oversaw the project, manuscript revisions and framing.
@@ A.9 Reconstruction Evaluations: Additional Information
A.10 UMAP Dimensionality Reduction
A.11 Pretraining with Less Subjects
@@
A.12 Enhanced Reconstructions using Empty Text Prompts
To investigate whether our generated text captions from brain activity were improving subsequent image generation, we conducted an additional ablation. We remade reconstructions for subject 1 using their fine-tuned MindEye2 model but used empty text prompts instead of the image captions predicted from brain activity. We observed modest improvements in high-level metrics due to the use of predicted image captions, as shown in Table  11.
Metrics Brain captions Empty prompt
PixCorr $\uparrow$ 0.235 0.235
SSIM $\uparrow$ 0.428 0.423
Alex(2) $\uparrow$ 88.02% 87.46%
Alex(5) $\uparrow$ 93.33% 93.22%
Incep $\uparrow$ 83.56% 82.67%
CLIP $\uparrow$ 80.75% 81.21%
Eff $\downarrow$ 0.798 0.810
SwAV $\downarrow$ 0.459 0.461
Table 11: Using image captions predicted from brain activity for refinement seems to modestly improve subsequent image reconstructions from fMRI activity compared to using empty text prompts.
A.13 Subject-specific semantically labeled UMAPs
Appendix Figure  10 shows subject-specific UMAP plots with labeled semantic clusters to visualize what categories of images were better or worse reconstructed for each subject. If a plot does not show dense clustering of images within the same semantic category, this indicates that reconstructions from that subject for this category were likely not of good quality. Broad semantic category labels were manually labeled and taken from a Notion report by Shirakawa (2023), which details his explorations of the Natural Scenes Dataset images.
Figure 10: Subject-specific UMAPs on the diffusion prior outputs from the 40-hour MindEye2 models help to visualize individual differences with regard to reconstuction quality for different categories of natural images.
A.14 ROI-Optimized Stimuli
Here we try to visualize the functional organization of the brain by feeding synthetic brain activity through pretrained MindEye2. Inspired by the ROI-optimal analyses of Ozcelik and VanRullen (2023) and the synthetic brain maximization approach of BrainDIVE (Luo et al., 2023a), we utilized four ROIs derived from population receptive field (pRF) experiments and four ROIs derived from functional localization (fLoc) experiments. These pRF and fLoc experiments were provided by the NSD dataset. The ROIs are as follows (region names following the terminology adopted in Allen et al. (2021)): V1 is the concatenation of V1 ventral (V1v) and V1 dorsal (V1d), and similarly for V2 and V3; V4 is the human V4 (hV4); the Face-ROI consists of the union of OFA, FFA-1, FFA-2, mTL-faces, and aTL-faces; the Word-ROI consists of OWFA, VWFA-1, VWFA-2, mfs-words, and mTL-words; the Place-ROI consists of OPA, PPA, and RSC; and the Body-ROI consists of EBA, FBA-1, FBA-2, and mTL-bodies.
@@ A.14 ROI-Optimized Stimuli
Figure 11: Unrefined reconstructions and decoded captions from synthetic fMRI activity. Voxels in the desired target brain region were set to a one while all other voxels were set to zero, and this synthetic brain data was fed through each subject’s MindEye2 model.
@@
Such results demonstrate the potential to directly visualize preferential stimuli for any desired region of interest; further functional specialization exploration could be performed using more sophisticated methods (c.f., Sarch et al. (2023); Luo et al. (2023a, b); Tang et al. (2022); Huth et al. (2016); Sucholutsky et al. (2023)).
A.15 Human Preference Experiments
@@
A.15.1 MindEye2 vs. MindEye2 (random)
@@
A.15.2 MindEye2 (refined) vs. MindEye2 (unrefined)
@@
A.15.3 MindEye2 (1-hour) vs. Brain Diffuser (1-hour)